# Fluctuation Theorem and Optimal Control of an Active Tracking Particle with Information Processing

### v1

@@ Abstract
We utilise a bipartite Bayesian network [42, 43] to describe the dynamics of the active particle, consisting of a physical subsystem $\mathcal{X}$ in charge of the particle locomotion as an ordinary run-and-tumble particle and a control (measurement - feedback) subsystem $\mathcal{C}$ in charge of the information-based regulation. With this treatment, the regulation of the whole process simply follows the classical Bellman description [44] of dynamic programming where the total optimal cost is the summation of all the costs at every time step, so it is sufficient to focus on the regulation process during a single time step, $t_{k}\rightarrow t_{k+1}$ with $0\leq k<\mathcal{T}$. For achieving regulatory purposes during any arbitrary time step, $t_{k}\rightarrow t_{k+1}$, we take the Maxwell-demon-like setup of the measurement - feedback process [45, 46, 47], and it is detailed as follows. At time $t_{k}$, the state of the particle is $(x_{k},n_{k})$, where $n_{k}=L,R$ with equal probability at equilibrium, i.e., $p_{\mathrm{eq}}^{0}(n_{k}=L,R)=1/2$. Then we measure its orientation denoted as $m_{k}$, and such measurements have an error, $0\leq\epsilon\leq 1$; in other words, we can define a conditional probability characterising the measurement: $p(m_{k}|n_{k})=1-\epsilon$ for $m_{k}=n_{k}$ and $p(m_{k}|n_{k})=\epsilon$ for $m_{k}\neq n_{k}$. If the measured orientation is pointing to the left, i.e., $m_{k}=L$, then we apply a magnetic field of the magnitude $B$ and the direction pointing to the right during a finite time interval $\Delta t$ to change its orientation; if the measured orientation is pointing to the right, i.e., $m_{k}=R$, then we do not apply the magnetic field and let the particle orientation freely evolve during $\Delta t$ by thermal fluctuation. We use $n_{k}^{\prime}$ to denote the particle orientation after the measurement and feedback control. The transition, $n_{k}\rightarrow n_{k}^{\prime}$, which is critical in the regulatory process, includes the nodes, $n_{k}$, $m_{k}$, and $n_{k}^{\prime}$ as shown in Fig. 1, and the transition can be described by the conditional probability, $p(n^{\prime}_{k}|n_{k},m_{k})$, with its explicit expressions shown in the table of the End Matter. For example, no magnetic field is applied in the case of ($n_{k}=L$, $m_{k}=R$), where the particle orientation will evolve freely during $\Delta t$, and we can obtain the probabilities of finding the particle with $n_{k}^{\prime}=L$ and $n_{k}^{\prime}=R$ as, $p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=L)+[1-p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=L)]\cdot\exp[-k_{0}\Delta t]$ and $p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=R)-[1-p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=L)]\cdot\exp[-k_{0}\Delta t]$, respectively, where the transition rate is $k_{0}=k_{0}^{L\rightarrow R}+k_{0}^{R\rightarrow L}=2k_{0}^{L\rightarrow R}$ with $k_{0}^{L\rightarrow R}$ and $k_{0}^{R\rightarrow L}$ as the transition rate from $n_{k}=L$ to $n_{k}^{\prime}=R$ and that from $n_{k}=R$ to $n_{k}^{\prime}=L$ without magnetic field, respectively. In following discussions, the notation $u_{p}=\exp[-k_{0}\Delta t]$ (related with unchanged population) is taken for simplicity. After the measurement - feedback process, we let the particle move along its regulated orientation with a constant step length, $\Delta x$, which is described by the process $(x_{k},n_{k}^{\prime})\rightarrow(x_{k+1},n_{k}^{\prime})$, and such dynamic process can be achieved in various active systems, e.g., by shining light on photocatalytical colloidal particles [48, 49] or applying electric field on metal-dielectric Janus particles [50, 51]. Then let the particle orientation relax to the state with its orientation distribution recovering to $p_{\mathrm{eq}}^{0}(n_{k+1}=L,R)=1/2$ at time $t_{k+1}$. For those familiar with bacteria swimming in the run-and-tumble mode, the above setup can be treated as a bacterium moving with a ‘brain’ which performs measurement and feedback control. By repeating the above process from $t_{0}$ to $t_{\mathcal{T}}$, the particle can arrive at its destination after $\mathcal{T}$ time steps, with its state as $(\mathcal{D}\Delta x,n_{\mathcal{T}})$.
@@
$\Delta s_{\mathcal{X}}^{k}=\frac{\ln p(n_{k})}{\ln p(n_{k+1})},$ (1)
@@
Figure 2: Mean first passage steps $\langle\mathcal{T}\rangle$ as a function of the magnetic field $\hat{B}$, for $\epsilon=0$ (red), 1/2 (blue) and 1 (brown). The black solid line is obtained by $\langle\mathcal{T}\rangle\simeq 2/[\hat{B}(1-u_{p})]$ for $\hat{B}\rightarrow 0$, and the round circles for different $\epsilon$ denote the values obtained by $\langle\mathcal{T}\rangle\simeq 2/(1+u_{p}-2\epsilon u_{p})$ for $\hat{B}\rightarrow\infty$. The parameter, $u_{p}=\exp[-k_{0}\Delta t]=0.9$ is taken for illustration.
@@
which depends on the process of measurement and feedback control through the probability, $p(n^{\prime}_{k}=L)$ [explicit expression shown in the table of the End Matter]. Note that $\langle\mathcal{T}\rangle$ diverges for unbiased random walk with $p(n^{\prime}_{k}=R)=1/2$. Apparently, the averaged first passage steps decreases with increasing magnetic field and decreasing measurement error, as shown in Fig. 2. We introduce the dimensionless field strength, $\hat{B}=B/B_{0}$ with $B_{0}=k_{B}T/m$ as the referenced magnetic field. For weak field controls, $\hat{B}\rightarrow 0$, the averaged first passage steps decreases with the increasing magnetic field in the form of $\langle\mathcal{T}\rangle\simeq 2/[\hat{B}(1-u_{p})]$, and for strong field controls, $\hat{B}\rightarrow\infty$, the averaged first passage steps only depends on the measurement error, in the form of $\langle\mathcal{T}\rangle\simeq 2/(1+u_{p}-2\epsilon u_{p})$.
@@ Acknowledgements.
$L,R,R$ $p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=R)-[1-p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=R)]\cdot e^{-k_{0}\Delta t}$
$R,R,L$ $p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=L)-[1-p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=L)]\cdot e^{-k_{0}\Delta t}$

### v3

@@ Abstract
We utilise a bipartite Bayesian network [42, 43] to describe the dynamics of the active particle, consisting of a physical subsystem $\mathcal{X}$ in charge of the particle locomotion as an ordinary run-and-tumble particle and a control (measurement - feedback) subsystem $\mathcal{C}$ in charge of the information-based regulation. With this treatment, the regulation of the whole process simply follows the classical Bellman description [44] of dynamic programming where the total optimal cost is the summation of all the costs at every time step, so it is sufficient to focus on the regulation process during a single time step, $t_{k}\rightarrow t_{k+1}$ with $0\leq k<\mathcal{T}$. For achieving regulatory purposes during any arbitrary time step, $t_{k}\rightarrow t_{k+1}$, we take the Maxwell-demon-like setup of the measurement - feedback process [45, 46, 47], and it is detailed as follows. At time $t_{k}$, the state of the particle is $(x_{k},n_{k})$, where $n_{k}=L,R$ with equal probability at equilibrium, i.e., $p_{\mathrm{eq}}^{0}(n_{k}=L,R)=1/2$. Then we measure its orientation denoted as $m_{k}$, and such measurements have an error, $0\leq\epsilon\leq 1$; in other words, we can define a conditional probability characterising the measurement: $p(m_{k}|n_{k})=1-\epsilon$ for $m_{k}=n_{k}$ and $p(m_{k}|n_{k})=\epsilon$ for $m_{k}\neq n_{k}$. If the measured orientation is pointing to the left, i.e., $m_{k}=L$, then we apply a magnetic field of the magnitude $B$ and the direction pointing to the right during a finite time interval $\Delta t$ to change its orientation; if the measured orientation is pointing to the right, i.e., $m_{k}=R$, then we do not apply the magnetic field and let the particle orientation freely evolve during $\Delta t$ by thermal fluctuation. We use $n_{k}^{\prime}$ to denote the particle orientation after the measurement and feedback control. The transition, $n_{k}\rightarrow n_{k}^{\prime}$, which is critical in the regulatory process, includes the nodes, $n_{k}$, $m_{k}$, and $n_{k}^{\prime}$ as shown in Fig. 1, and the transition can be described by the conditional probability, $p(n^{\prime}_{k}|n_{k},m_{k})$, with its explicit expressions shown in the table of the Appendix. For example, no magnetic field is applied in the case of ($n_{k}=L$, $m_{k}=R$), where the particle orientation will evolve freely during $\Delta t$, and we can obtain the probabilities of finding the particle with $n_{k}^{\prime}=L$ and $n_{k}^{\prime}=R$ as, $p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=L)+[1-p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=L)]\cdot\exp[-k_{0}\Delta t]$ and $p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=R)-[1-p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=L)]\cdot\exp[-k_{0}\Delta t]$, respectively, where the transition rate is $k_{0}=k_{0}^{L\rightarrow R}+k_{0}^{R\rightarrow L}=2k_{0}^{L\rightarrow R}$ with $k_{0}^{L\rightarrow R}$ and $k_{0}^{R\rightarrow L}$ as the transition rate from $n_{k}=L$ to $n_{k}^{\prime}=R$ and that from $n_{k}=R$ to $n_{k}^{\prime}=L$ without magnetic field, respectively. In following discussions, the notation $u_{p}=\exp[-k_{0}\Delta t]$ (related with unchanged population) is taken for simplicity. After the measurement - feedback process, we let the particle move along its regulated orientation with a constant step length, $\Delta x$, which is described by the process $(x_{k},n_{k}^{\prime})\rightarrow(x_{k+1},n_{k}^{\prime})$, and such dynamic process can be achieved in various active systems, e.g., by shining light on photocatalytical colloidal particles [48, 49] or applying electric field on metal-dielectric Janus particles [50, 51]. Then let the particle orientation relax to the state with its orientation distribution recovering to $p_{\mathrm{eq}}^{0}(n_{k+1}=L,R)=1/2$ at time $t_{k+1}$. For those familiar with bacteria swimming in the run-and-tumble mode, the above setup can be treated as a bacterium moving with a ‘brain’ which performs measurement and feedback control. By repeating the above process from $t_{0}$ to $t_{\mathcal{T}}$, the particle can arrive at its destination after $\mathcal{T}$ time steps, with its state as $(\mathcal{D}\Delta x,n_{\mathcal{T}})$.
@@
$\Delta s_{\mathcal{X}}^{k}=\ln\frac{p(n_{k})}{p(n_{k+1})},$ (1)
@@
Figure 2: Mean first passage steps $\langle\mathcal{T}\rangle$ as a function of the magnetic field $\hat{B}$, for $\epsilon=0$ (red), 1/2 (blue) and 1 (brown). The black solid line is obtained by $\langle\mathcal{T}\rangle/\mathcal{D}\simeq 2/[\hat{B}(1-u_{p})]$ for $\hat{B}\rightarrow 0$, and the round circles for different $\epsilon$ denote the values obtained by $\langle\mathcal{T}\rangle/\mathcal{D}\simeq 2/(1+u_{p}-2\epsilon u_{p})$ for $\hat{B}\rightarrow\infty$. The parameter, $u_{p}=\exp[-k_{0}\Delta t]=0.9$ is taken for illustration.
@@
which depends on the process of measurement and feedback control through the probability, $p(n^{\prime}_{k}=L)$ [explicit expression shown in the table of the Appendix]. Note that $\langle\mathcal{T}\rangle$ diverges for unbiased random walk with $p(n^{\prime}_{k}=R)=1/2$. Apparently, the averaged first passage steps decreases with increasing magnetic field and decreasing measurement error, as shown in Fig. 2. We introduce the dimensionless field strength, $\hat{B}=B/B_{0}$ with $B_{0}=k_{B}T/m$ as the referenced magnetic field. For weak field controls, $\hat{B}\rightarrow 0$, the averaged first passage steps decreases with the increasing magnetic field in the form of $\langle\mathcal{T}\rangle/\mathcal{D}\simeq 2/[\hat{B}(1-u_{p})]$, and for strong field controls, $\hat{B}\rightarrow\infty$, the averaged first passage steps only depends on the measurement error, in the form of $\langle\mathcal{T}\rangle/\mathcal{D}\simeq 2/(1+u_{p}-2\epsilon u_{p})$.
@@ Acknowledgements.
$L,R,R$ $p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=R)-[1-p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=L)]\cdot e^{-k_{0}\Delta t}$
$R,R,L$ $p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=L)-[1-p_{\mathrm{eq}}^{0}(n_{k}^{\prime}=R)]\cdot e^{-k_{0}\Delta t}$